Operations Report – February 18th

Crew 189 Operations Report 18Feb2018

SOL: 0

Name of person filing report: Gabriel PAYEN

Non-nominal systems: Pipe leak in low deck, water heater, two thermometers in the Greenhab.

Notes on non-nominal systems: –

Generator (hours run): 14 hours during the night (7 PM to 9 AM)

Solar— SOC % (Before generator is run at night) : 91%

Diesel – ?

Propane – 46.2 psi.

Ethanol Free Gasoline (5 Gallon containers for ATV) – 7 gallons (estimation based the total height and capacity of the container)

Water (trailer) – 0 gallons.

Water (static) – 400 gallons (estimation based the total height and capacity of the tank)

Trailer to Static Pump used – No

Water (loft) – Static to Loft Pump used – Yes

Water Meter (loft) : 0.44, measurement at 7 PM

Toilet tank emptied: No

ATV’s Used: All ATVs used except the one which doesn’t work (I forgot to check its number)

Oil Added : No

ATV Fuel Used: Gals

# Hours the ATVs were Used today: 20min

Notes on ATVs: –

Deimos rover used: Yes

Hours: 108.7

Beginning charge: ? (not checked)

Ending charge: 98%

Currently charging: No

Spirit rover used: No

Hours: 20.2

Beginning charge: –

Ending charge: –

Currently charging: Yes (Atilia may have unplugged Curiosity to plug in Spirit)

Opportunity rover used: Yes

Hours: 15.3

Beginning charge: ? (not checked)

Ending charge: 93% (100% after being charged)

Currently charging: No

Curiosity rover used: No

Hours: 12.7

Beginning charge: 100%

Ending charge: 100%

Currently charging: No

HabCar used and why, where : Not used

General notes and comments: Last checks today with Shannon, some things to fix (see Non-nominal systems). Jérémy tried to run his aerosol counter device (LOAC) in the Hab but it doesn’t work. It turns on but doesn’t begin counting. The crew will try to fix it tomorrow.

Summary of internet: –

Summary of suits and radios: All air pipes had been tested and only one doesn’t work (backpack 3 – left one). All radios had been tested and work

Summary of Hab operations: –

Summary of GreenHab operations: –

Summary of ScienceDome operations: –

Summary of RAMM operations: –

Summary of health and safety issues: –

Questions, concerns and requests to Mission Support:


The GreenHab officer would need a new thermometer.

-Pipe leaks

A bucket is beneath the pipe and receives a liquid which seems to be bad smelling water. Tape would stop the flow, but it only seems to be a temporary solution. Tomorrow I’ll try to find where the leak comes from.

-Water heater

Shannon seems to have fixed it before the very beginning of the SIM. We’ll see it when a crew member will take the first shower of the mission.


Some checks have been forgotten (beginning charge of the rovers, numbers of ATVs used…). Sorry for that, I’ll try not to do theses mistakes anymore.

Copyright © The Mars Society. All rights reserved. | Main Site